OpenStudy (anonymous):

sec^4x = (1+tan^2x) (sec^2x) Please Help Me

OpenStudy (anonymous):

with? Trig identities?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1+tan^2x is sec^2x

OpenStudy (anonymous):

Understood that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is sec^4x

OpenStudy (anonymous):

is it 1+tan^4x

OpenStudy (anonymous):

since 1+tan^2x=sec^2x sec^2x x sec^2x= sec^4x

OpenStudy (irishboy123):

\( \large sec^4x = (1+tan^2x) (sec^2x) \)  \( \large \frac{h^4}{a^4} = (1 + \frac{o^2}{a^2} )(\frac{h^2}{a^2})\) \(\large h^4 = (a^2 +o^2)h^2 = h^2.h^2\) Q.E.D.